A 0.1 kg mass is suspended from a wire of negligible mass.The length of the wire is 1m and it`s cross sectional area is 4.9×10 power-7 sq metre.if the mass is pulled little in the vertically downward direction and released,it performs SHM of angular frequency 140 rad/sec.if the Young`s modulus of the material of the wire is a×10power9 N/ sq metre,value of a is

Grade:12

1 Answers

Arun
6 years ago
 
We know that 8 = vK/m ……..(i)
Here Y = FL/Al ? F = (YA/L)l
Comparing the above equations with F = kl we get
K= (YA/L) ………….(ii)
From (i) & (ii), ? = v YA/ml
? 140 = vn*109 * 4.9 * 10-7/0.1 * 1
? n =4